Bob is 4 times greater than Susan. Dakota is 3 years younger then Susane. The sum of bob, Susan, and Dakota ages is 93. What is Susanes age
Answer by Cromlix(4381)

Susan's age = x
Bob's age is 4x
Dakota's age = x - 3
Total ages = 93
x + 4x + (x - 3) = 93
6x - 3 = 93
6x = 96
x = 16
Susan's age = 16
Bob's age = 64
Dakota's age = 13.
Hope this helps.
:-)